# The Repeat Offenders: Characterizing and Predicting Extremely Bug-Prone Source Methods

### v1

@@ Abstract.
Identifying the small subset of source code that repeatedly attracts bugs is critical for reducing long-term maintenance effort. We define ExtremelyBuggy methods as those involved in more than one bug fix and present the first large-scale study of their prevalence, characteristics, and predictability. Using a dataset of over 1.25 million methods from 98 open-source Java projects, we find that ExtremelyBuggy methods constitute only a tiny fraction of all methods, yet frequently account for a disproportionately large share of bugs. At their inception, these methods are significantly larger, more complex, less readable, and less maintainable than both singly-buggy and non-buggy methods. However, despite these measurable differences, a comprehensive evaluation of five machine learning models shows that early prediction of ExtremelyBuggy methods remains highly unreliable due to data imbalance, project heterogeneity, and the fact that many bugs emerge through subsequent evolution rather than initial implementation. To complement these quantitative findings, we conduct a thematic analysis of 265 ExtremelyBuggy methods, revealing recurring visual issues (e.g., confusing control flow, poor readability), contextual roles (e.g., core logic, data transformation, external resource handling), and common defect patterns (e.g., faulty conditionals, fragile error handling, misuse of variables). These results highlight the need for richer, evolution-aware representations of code and provide actionable insights for practitioners seeking to prioritize high-risk methods early in the development lifecycle.
ExtremelyBuggy, software bug, bug-proneness, bug prediction, software maintenance, code metrics.
$c$s: Software and its engineering Maintaining software
@@ 1. Introduction
Software maintenance is one of the most expensive phases of the development lifecycle (borstler_role_2016). A major contributor to this cost is the effort required to identify and correct software bugs, which alone can account for 50-70% of the total development costs (wang2023bugpre; chowdhury_method-level_2024). To no surprise, building models for accurate bug prediction is referred to as the prince of empirical software engineering research (lanza2016tragedy) with researchers developing bug prediction models in an attempt to take early actions (chowdhury_method-level_2024; giger_method-level_2012; pascarella_performance_2020; rahman2017relationships; shivaji2012reducing; wang2023bugpre; wattanakriengkrai2020predicting). The assumption being that if bug-prone code components can be identified early in the development lifecycle or prevented altogether, future maintenance costs will be significantly reduced (ibm).
Prior bug prediction models predominantly focused on the class or file levels (basili_validation_1996; gil_correlation_2017; zhou_ability_2010). Unfortunately, these models remain underutilized in practice (tantithamthavorn2018experience), at least partly because practitioners find it challenging to locate bugs at such coarse levels of granularity (chowdhury_method-level_2024; grund_codeshovel_2021; pascarella_performance_2020; lanza2016tragedy; shihab_industrial_2012; mondal_investigating_2019). Practical use would require developers to sift through entire classes or files without an indication of what characteristics the model considered bug-prone. Consequently, a significant amount of recent research has focused on bug prediction at the method level granularity (mo2022exploratory; giger_method-level_2012; chowdhury_method-level_2024; Shippey; ferenc_automatically_2020; pascarella_performance_2020). However, these models treat all buggy methods equally, without distinguishing between methods that were fixed once and those that required multiple bug fixes.
@@
Contribution 1: We found that 0.04-6.63% of methods are ExtremelyBuggy. However, this small proportion of methods can often account for a large number of bug fixes in a project. This implies that identifying this small fraction of methods early will significantly reduce future maintenance burden.
@@
Contribution 2: We found that, at their inception, ExtremelyBuggy methods are significantly larger, less readable, and have lower maintainability scores compared to both Buggy and NotBuggy methods. This is encouraging, as it suggests that these methods exhibit distinguishable code quality that could be leveraged by machine learning models for early prediction. That led us to our next research question.
@@
Contribution 4: The futility in predicting with machine learning models led us to manually investigate the characteristics of the ExtremelyBuggy methods so that we can provide actionable guidance for the practitioners. As such, we conducted a thematic analysis on a curated dataset of 287 ExtremelyBuggy methods. In general, we focused on three different aspects while applying thematic analysis: visual issues, context of a method, and bug-fix type. Our analysis reveals some common themes that exist within these 287 methods. For example, we found that methods that deal with the core logic and algorithms tend to become ExtremelyBuggy.
To help replication and extension in future method-level bug prediction research, we share this dataset publicly.  [^1] https://github.com/SQMLab/ExtremelyBuggyPublicData/tree/main/Dataset
@@ 2.1. Software Maintenance and Bug Prediction
Software maintenance is expensive (borstler_role_2016) and the bug-proneness of code components has been identified as one of the largest indicators of future maintenance burdens (chowdhury_good_2025). As such, there has been an enormous effort by the research community over the last forty years (lenarduzzi2017analyzing) to predict bug-prone code components so that early actions can be taken (d2010extensive; chowdhury_good_2025). Despite the great amount of research dedicated to predicting bugs, bug prediction models have yet to see widespread industry use (lanza2016tragedy; tantithamthavorn2018experience). This is due, at least in part, to the fact that prior studies have predominantly focused on bug prediction at the file or class level (basili_validation_1996; gil_correlation_2017; zhou_ability_2010), a granularity that practitioners often consider too coarse for practical use (mondal_investigating_2019; pascarella_performance_2020; grund_codeshovel_2021; chowdhury_method-level_2024).
These concerns have led to a recent shift in bug prediction research towards the method level (ferenc_automatically_2020; giger_method-level_2012; hata2012bug) granularity. However, multiple recent studies (chowdhury_method-level_2024; pascarella_performance_2020) have shown that the high prediction accuracy [^2] Unless otherwise stated, accuracy in this paper refers to precision, recall, and F1-score. reported by these studies is often inflated due to unrealistic evaluation scenarios, and, when tested on realistic scenarios, the accuracy of these models drops significantly. Additionally, prior studies have not differentiated between methods associated with a single bug fix and those that are repeatedly involved in bug-fixing. This distinction is critical, as methods that frequently require maintenance—often referred to as maintenance hotspots—should be prioritized and allocated additional resources.
Motivated by this important gap in prior research, this paper focuses exclusively on source code methods that have been associated with bugs more than once. We investigate whether these methods exhibit distinct characteristics and whether they can be identified early in their lifecycle.
@@ 2.2. Code Metrics
For bug prediction models to work, they need some quantifiable metrics that characterize the source code to use as inputs to the models. These metrics often come in the form of product metrics (ferenc_automatically_2020; mo2022exploratory; pascarella_performance_2020; chowdhury_revisiting_2022) and process metrics (graves2002predicting; mo2022exploratory; rahman2013and). Process metrics are typically derived from the change history of a code component and have been shown to be effective predictors of bug-prone components (giger_method-level_2012; rahman2013and). However, such metrics are not available immediately after a piece of code is written and therefore cannot support the early identification of maintenance-prone components. Waiting for sufficient change history to accumulate may allow these methods to cause significant issues, leading to high maintenance costs and potentially eroding stakeholder confidence.
Product metrics, on the other hand, are often readily available and easy to compute at the inception of a code component. As a result, building maintenance prediction models using only product metrics has been a long-standing goal in software engineering research (romano_using_2011; pascarella_performance_2020; chowdhury_good_2025; nunez2017source). Despite their extensive usage in software engineering research, some studies have found that product metrics, apart from size, are not helpful for building models that predict future maintenance burdens (gil_correlation_2017; el2001confounding; sjoberg2012quantifying). However, these studies with negative results were conducted at the class/file level, whereas studies that focused on method-level granularity have found product metrics as useful predictors of maintenance (landman2014empirical; chowdhury_method-level_2024; chowdhury_evidence_2025; chowdhury_revisiting_2022).
Motivated by the need for early prediction and the practicality of method-level source code metrics, we focus exclusively on these metrics to characterize and predict ExtremelyBuggy methods.
@@ 3. Methodology
In this section, we outline the overall methodology, including project selection, change history tracking, and the identification of bug-proneness indicators. We also detail the data preprocessing steps and the selection of code metrics. To improve readability, methodologies specific to individual research questions are presented alongside each corresponding research question.
@@ 3.1. Dataset
A substantial portion of our dataset originates from the work of Chowdhury et al. (chowdhury_method-level_2024), which comprises 774,051 Java methods drawn from 49 open-source projects. Each method in this dataset is accompanied by its complete change history and is enriched with metadata such as commit messages, timestamps, authorship information, and change types. The dataset additionally includes code diffs for all historical modifications, binary labels indicating whether each change corresponds to a bug fix, and the number of methods modified within the associated commit. These 49 projects have been widely utilized, either fully or partially, in prior method-level software engineering studies (ray_naturalness_2016; gil_correlation_2017; Palomba:2017; spadini_relation_2018; grund_codeshovel_2021; chowdhury_evidence_2025; chowdhury_method-level_2024; chowdhury_good_2025). Building upon this foundation, we extend the dataset by collecting additional open-source projects and extracting their method-level change histories.
@@ 3.1.1. Project Selection
We employed the GitHub REST API [^3] https://docs.github.com/en/rest and the PyGitHub library [^4] https://github.com/PyGithub/PyGithub to collect open-source projects. To ensure the quality and reliability of our dataset, we followed recommended best practices for mining GitHub data (kalliamvakou2014promises). We began by retrieving the 1000 most-starred GitHub repositories that satisfied several additional criteria. First, to reflect contemporary industry practices, we restricted our search to repositories created within the past 15 years [^5] Created after January 1st, 2010 and excluded repositories that had been inactive for more than one year [^6] No commits since May 13, 2024. Second, to ensure representativeness of real-world Java projects, we excluded templates, archived repositories, and forks by specifying appropriate parameters in our GitHub API queries. Third, we required repositories to contain at least $2000$ commits to guarantee substantial change history. We further filtered out repositories whose codebase consisted of less than $95\%$ Java, thereby ensuring the availability of a sufficiently large number of Java methods for extraction.
Following the automated filtering process, we manually reviewed all remaining repositories to verify adherence to the above criteria and to confirm that their descriptions were written in English. Ensuring English descriptions facilitates the identification of bug-fixing commits using English-language keywords and supports subsequent manual qualitative analysis of commit messages. Based on this review, we selected 49 additional projects. Combined with the 49 projects from Chowdhury et al. (chowdhury_method-level_2024), our final dataset comprises 98 projects in total. The complete list of filtered projects is included in our repository  [^7] https://github.com/SQMLab/ExtremelyBuggyPublicData/tree/main/AdditionalFigures/RepositoryList.md.
@@ 3.1.2. Extracting Method-level History
To quantify the bug-proneness of a method, it is necessary to determine how frequently that method has been involved in bug-fixing activities, which in turn requires capturing its complete change history. To obtain this historical information, we employed the CodeShovel tool (grund_codeshovel_2021). CodeShovel has demonstrated high accuracy in identifying method-level change histories and is robust to complex code transformations, including method relocations across files (grund_codeshovel_2021). Its reliability has also been independently evaluated by industry practitioners, who reported similarly strong accuracy (grund_codeshovel_2021). Although a more recent tool, CodeTracker (jodavi_accurate_2022), has been shown to achieve higher accuracy than CodeShovel, its substantially slower performance renders it impractical for tracing the histories of large numbers of methods. Moreover, CodeTracker has not undergone independent validation by industry professionals.
@@
To support large-scale and reliable analysis, we used CodeShovel to generate structured JSON outputs for approximately 1.25 million methods across 98 projects, capturing their complete method-level evolution across commits. These outputs provide a robust foundation from which we extract both code metrics and change metrics for every method revision included in our dataset.

### v2

# arXiv:2511.22726v2[cs.SE] 27 Feb 2026

@@ Abstract.
Bug prediction has long been considered the “prince” of empirical software engineering research, and accordingly, a substantial body of work has focused on predicting bugs to enable early preventive actions. However, most existing studies operate at the class or file level, which practitioners have found to be of limited practical value. As a result, method-level bug prediction has gained increasing attention in recent years. Despite this shift, current method-level prediction models typically treat all buggy methods as equally fault-prone, regardless of whether a method has been associated with a bug once or repeatedly. We argue that methods involved in bugs multiple times—hereafter referred to as ExtremelyBuggy methods—are more harmful than methods that are buggy only once. In this study, we investigate the prevalence of ExtremelyBuggy methods, analyze their code quality metrics, and assess whether they can be predicted at the time of their introduction. In addition, we conduct a thematic analysis of 287 ExtremelyBuggy methods to gain deeper insights into their characteristics.
Using a dataset of over 1.25 million methods extracted from 98 open-source Java projects, we find that only a small proportion of methods can be classified as ExtremelyBuggy, yet these methods account for a disproportionately large share of bugs within a project. Although we observe statistically significant differences between ExtremelyBuggy and other methods, ExtremelyBuggy methods remain difficult to predict at their inception. Nevertheless, our manual analysis reveals recurring characteristics among these methods. These findings can help practitioners avoid harmful patterns in practice and guide future research toward developing features and models that better capture the unique properties of such methods.
software bug, bug-proneness, bug prediction, software maintenance, code metrics.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Software maintenance is one of the most expensive phases of the development lifecycle (Börstler and Paech, 2016). A major contributor to this cost is the effort required to identify and correct software bugs, which alone can account for 50-70% of the total development costs (Wang et al., 2023; Chowdhury et al., 2024). To no surprise, building models for accurate bug prediction is referred to as the prince of empirical software engineering research (Lanza et al., 2016) with researchers developing bug prediction models in an attempt to take early actions (Chowdhury et al., 2024; Giger et al., 2012; Pascarella et al., 2020; Rahman and Roy, 2017; Shivaji et al., 2012; Wang et al., 2023; Wattanakriengkrai et al., 2020). The assumption is that if bug-prone code components can be identified early in the development lifecycle or prevented altogether, future maintenance costs will be significantly reduced (ibm).
Prior bug prediction models predominantly focused on the class or file levels (Basili et al., 1996; Gil and Lalouche, 2017; Zhou et al., 2010). Unfortunately, these models remain underutilized in practice (Tantithamthavorn and Hassan, 2018), at least partly because practitioners find it challenging to locate bugs at such coarse levels of granularity (Chowdhury et al., 2024; Grund et al., 2021; Pascarella et al., 2020; Lanza et al., 2016; Shihab et al., 2012; Mondal et al., 2019). Practical use would require developers to inspect the entire flagged class or file to identify the specific location of the bug—a laborious and time-consuming task. Consequently, a significant amount of recent research has focused on bug prediction at the method-level granularity (Mo et al., 2022; Giger et al., 2012; Chowdhury et al., 2024; Shippey et al., 2016; Ferenc et al., 2020; Pascarella et al., 2020). However, these models treat all buggy methods equally, without distinguishing between methods that were fixed once and those that required multiple bug fixes.
@@
Contribution 1: We found that 0.04-6.63% of methods are ExtremelyBuggy. However, this small proportion of methods can often account for a large number of bugs in a project (even more than 90% of the bugs in some projects). This implies that identifying this small fraction of methods early will significantly reduce future maintenance burden.
@@
Contribution 2: We found that, at their inception, ExtremelyBuggy methods are significantly larger, less readable, and have lower maintainability scores compared to both Buggy and NotBuggy methods. This is encouraging, as it suggests that these methods exhibit distinguishable code quality that could be leveraged by machine learning models for early prediction. This observation led us to our next research question.
@@
Contribution 4: The futility in predicting with machine learning models led us to manually investigate the characteristics of the ExtremelyBuggy methods so that we can provide actionable guidance for the practitioners. As such, we conducted a thematic analysis (Fereday and Muir-Cochrane, 2006) on a curated dataset of 287 ExtremelyBuggy methods. In general, we focused on three different dimensions while applying thematic analysis: visual issues, context of a method, and bug-fix type. Our analysis reveals some common themes that exist within these 287 methods. For example, we found that methods exhibiting self-admitted technical debt, implementing core logic or algorithms, or containing excessive exception handling or logging tend to become ExtremelyBuggy.
To help replication and extension in future method-level bug prediction research, we share our dataset publicly. [^1] https://github.com/SQMLab/ExtremelyBuggyPublicData/tree/main/Dataset
@@ 2.1. Software Maintenance and Bug Prediction
Software maintenance is expensive (Börstler and Paech, 2016) and the bug-proneness of code components has been identified as one of the largest indicators of future maintenance burdens (Chowdhury, 2025). In fact, reliably predicting software bugs is considered one of the holy grails (D’Ambros et al., 2010b) and the prince of empirical software engineering research (Lanza et al., 2016). As such, there has been an enormous effort  (Bangash et al., 2020; Tantithamthavorn and Hassan, 2018; D’Ambros et al., 2010a; Chowdhury, 2025; Rahman and Roy, 2017; Shivaji et al., 2012; Wang et al., 2023; Wattanakriengkrai et al., 2020; He et al., 2015) by the research community over the last forty years (Lenarduzzi et al., 2017) to predict bug-prone code components so that early actions can be taken. Despite the great amount of research dedicated to predicting bugs, bug prediction models have yet to see widespread industry use (Lanza et al., 2016; Tantithamthavorn and Hassan, 2018). This is due, at least in part, to the fact that prior studies have predominantly focused on bug prediction at the file or class level (Basili et al., 1996; Gil and Lalouche, 2017; Zhou et al., 2010), a granularity that practitioners often consider too coarse for practical use (Mondal et al., 2019; Pascarella et al., 2020; Grund et al., 2021; Chowdhury et al., 2024).
Although the most helpful models for practitioners would be to predict bugs at the line level, such models, however, do not offer high precision (Wattanakriengkrai et al., 2020), as many lines can be similar just by chance (Grund et al., 2021). Considering these factors, method-level bug prediction gained traction in recent years (Ferenc et al., 2020; Giger et al., 2012; Hata et al., 2012; Chowdhury et al., 2024; Pascarella et al., 2020; Shippey et al., 2016). For example, Ferenc et al. (Ferenc et al., 2020) showed that accurate prediction models can be built by using different product and process metrics. Similar results were reported by Giger et al. (Giger et al., 2012) and Hata et al. (Hata et al., 2012). Unfortunately, multiple recent studies (Chowdhury et al., 2024; Pascarella et al., 2020) have shown that the high prediction accuracy reported by these studies is often inflated due to unrealistic evaluation scenarios. These studies have used k-fold cross-validation for model evaluation, which is prone to leakage between train and test data due to the time-sensitive nature of bug prediction data. And when tested on realistic scenarios, the accuracy of these models drops significantly (Chowdhury et al., 2024; Pascarella et al., 2020).
Additionally, prior studies have not distinguished between methods involved in a single bug fix and those repeatedly involved in bug fixing. This distinction is critical because methods that frequently require maintenance—often referred to as maintenance hotspots—should be prioritized and allocated additional resources. Moreover, such methods may be more distinctive than ordinary bug-prone methods, a property that could be leveraged by prediction models to improve their accuracy. Motivated by this important gap in prior research, this paper focuses exclusively on source code methods that have been associated with bugs more than once. We investigate whether these methods exhibit distinct characteristics and whether they can be identified early in their lifecycle.
@@ 2.2. Code Metrics
For bug prediction models to work, they need some quantifiable metrics that characterize the source code to use as inputs to the models. These metrics often come in the form of product metrics (Ferenc et al., 2020; Mo et al., 2022; Pascarella et al., 2020; Chowdhury et al., 2022b) and process metrics (Graves et al., 2002; Mo et al., 2022; rahman2013and). Process metrics are typically derived from the change history of a code component and have been shown to be effective predictors of bug-prone components (Giger et al., 2012; rahman2013and). However, such metrics are not available immediately after a piece of code is written and therefore cannot support the early identification of maintenance-prone components. Waiting for sufficient change history to accumulate may allow these methods to cause significant issues, leading to high maintenance costs and potentially eroding stakeholder confidence.
Product metrics, on the other hand, are readily available and easy to compute at the inception of a code component. As a result, building maintenance prediction models using only product metrics has been a long-standing goal in software engineering research (Romano and Pinzger, 2011; Pascarella et al., 2020; Chowdhury, 2025; Nuñez-Varela et al., 2017). Despite their extensive usage in software engineering research, some studies have found that product metrics, apart from size, are not helpful for building models that predict future maintenance burdens (Gil and Lalouche, 2017; El Emam et al., 2001; Sjøberg et al., 2012). However, these studies with negative results were conducted at the class/file level, whereas studies that focused on method-level granularity have found product metrics as useful predictors of maintenance (Landman et al., 2014; Chowdhury et al., 2024, 2025, 2022b).
Motivated by the need for early prediction and the practicality of method-level source code metrics, we focus exclusively on these metrics to characterize and predict ExtremelyBuggy methods. Our goal is to predict these harmful methods as soon as they are pushed to a system.
@@ 3. Methodology
Similar to previous method-level bug prediction studies (Pascarella et al., 2020; Chowdhury et al., 2024; Mo et al., 2022; Shippey et al., 2016), we focused on Java projects only. This also mitigates threats associated with code metric–based studies that rely on projects written in multiple programming languages (Zhang et al., 2013). In this section, we outline the overall methodology, as presented in Figure 1, including project selection, change history tracking, and the identification of bug-proneness indicators. We also detail the data preprocessing steps and the selection of code metrics. To improve readability, methodologies specific to individual research questions are presented alongside each corresponding research question.
Figure 1. Methodology for collecting, preprocessing, labeling and analyzing our dataset.
@@ 3.1. Dataset
We need to build a dataset of methods along with their change histories, so that we can label each method’s bug-proneness by identifying bug- and fix-related keywords in all the commit messages associated with that method. A substantial portion of our dataset originates from the work of Chowdhury et al. (Chowdhury et al., 2024), which comprises 774,051 Java methods drawn from 49 open-source projects. Each method in this dataset is accompanied by its complete change history and is enriched with metadata such as commit messages, timestamps, authorship information, and change types. The dataset additionally includes code diffs for all historical modifications, binary labels indicating whether each change corresponds to a bug fix, and the number of methods modified within the associated commit. These 49 projects have been widely utilized, either fully or partially, in prior method-level software engineering studies (Ray et al., 2016; Gil and Lalouche, 2017; Palomba et al., 2017; Spadini et al., 2018; Grund et al., 2021; Chowdhury et al., 2025, 2024; Chowdhury, 2025). Since ExtremelyBuggy methods are even rarer than buggy methods (as shown later), we aimed to expand the dataset of Chowdhury et al. Therefore, we collected additional open-source projects and extracted their method-level change histories.
@@ 3.1.1. Project Selection
We employed the GitHub REST API [^2] https://docs.github.com/en/rest and the PyGitHub library [^3] https://github.com/PyGithub/PyGithub to collect open-source projects. To ensure the quality and reliability of our dataset, we followed recommended best practices for mining GitHub data (Kalliamvakou et al., 2014). We began by retrieving the 1000 most-starred GitHub repositories that satisfied several additional criteria. First, to reflect contemporary industry practices, we restricted our search to repositories created within the past 15 years [^4] Created after January 1st, 2010 and excluded repositories that had been inactive for more than one year. [^5] No commits since May 13, 2024 Second, to ensure representativeness of real-world Java projects, we excluded templates, archived repositories, and forks by specifying appropriate parameters in our GitHub API queries. Third, we required repositories to contain at least $2000$ commits to guarantee substantial change history. We further filtered out repositories whose codebase consisted of less than $95\%$ Java, thereby ensuring the availability of a sufficiently large number of Java methods for extraction.
Following the automated filtering process, we manually reviewed all remaining repositories to verify adherence to the above criteria and to confirm that their descriptions were written in English. Ensuring English descriptions facilitates the identification of bug-fixing commits using English-language keywords and supports subsequent manual qualitative analysis of commit messages. After this review process, we selected 49 additional projects from the available projects. Combined with the 49 projects from Chowdhury et al. (Chowdhury et al., 2024), our final dataset comprises 98 projects in total. The complete list of filtered projects is included in our repository. [^6] https://github.com/SQMLab/ExtremelyBuggyPublicData/tree/main/AdditionalFigures/RepositoryList.md
@@ 3.1.2. Extracting Method-level History
To quantify the bug-proneness of a method, it is necessary to determine how frequently that method has been involved in bug-fixing activities, which in turn requires capturing its complete change history. To obtain this historical information, similar to Chowdhury et al. (Chowdhury et al., 2024), we employed the CodeShovel tool (Grund et al., 2021). CodeShovel has demonstrated high accuracy in identifying method-level change histories and is robust to complex code transformations, including method relocations across files (Grund et al., 2021). Its reliability has also been independently evaluated by industry practitioners, who reported similarly strong accuracy (Grund et al., 2021). Although a more recent tool, CodeTracker (Jodavi and Tsantalis, 2022), has been shown to achieve higher accuracy than CodeShovel, its substantially slower performance made it challenging to use it for tracing the histories of the large number of methods we have in our dataset. Moreover, CodeTracker has not undergone independent validation by industry professionals.
@@
To support large-scale and reliable analysis, we gathered structured JSON outputs for approximately 1.25 million methods across 98 projects, capturing their complete method-level evolution across commits. These outputs provide a robust foundation from which we extract both code metrics and change metrics for every method revision included in our dataset. To the best of our knowledge, no prior bug prediction study has included such a large number of projects and methods. We believe this dataset can serve as a valuable resource for future research on method-level bug prediction.
3.2. Code Metrics
Code metrics have traditionally served as indicators of quality and maintainability of code (Gil and Lalouche, 2017; Pascarella et al., 2020; Chowdhury et al., 2024). Encouraged by previous research in method-level code metrics (Mo et al., 2022; Pascarella et al., 2020; Chowdhury et al., 2025), we used 14 code metrics to determine if the code quality indicators of the ExtremelyBuggy methods are different than others. These metrics were calculated using the same tool used in  (Chowdhury et al., 2025, 2024, 2022b). The selected metrics are described as follows.
Size. Size is one of the simplest metrics, yet it is considered one of the most important with regard to software quality and maintenance (Gil and Lalouche, 2017; El Emam et al., 2001; Chowdhury et al., 2022a). We used the number of source lines of code without comments and blank lines as our size measurement (referred to as SLOCStandard) (Chowdhury et al., 2024; Landman et al., 2014; Ralph and Tempero, 2018; Chowdhury et al., 2025).
Readability. The ability to read and understand existing code is an important part of software maintenance (Scalabrino et al., 2016). Code readability can be broadly defined by how easy it is for a developer to understand the structure and flow of source code (Buse and Weimer, 2009; Posnett et al., 2011; Börstler and Paech, 2016; Johnson et al., 2019). Developers can more easily understand and therefore modify readable code with reduced risk of issues. We used two distinct readability metrics in this paper: Buse et al. (Buse and Weimer, 2009) (Readability) and Posnett et al. (Posnett et al., 2011) (SimpleReadability). Both provide a value for each method in the range $[0,1]$.
Complexity and Testability. We used the McCabe metric (McCabe, 1976) to analyze the complexity of methods. This metric represents the number of independent paths in a method; the greater the number of paths, the more challenging it becomes to test the method thoroughly. In addition, we used the number of control variables (NVAR), the number of comparisons (NCOMP), and the McClure metric (McClure, 1978), which is calculated as the sum of NVAR and NCOMP. These metrics offer additional insights into method complexity that are not captured by the McCabe metric. Hindle et al. (Hindle et al., 2008) proposed the proxy indentation metric (IndentSTD (Chowdhury et al., 2022b)), which performs similarly to the McCabe metric in capturing code complexity but, unlike McCabe, does not require a language-specific parser. We included IndentSTD in our metric set as well. Finally, we used the MaximumBlockDepth metric, which measures the maximum nested block depth, as deep nesting within a method can further complicate testing.
Maintainability. Maintainability Index (MI) (Oman and Hagemeister, 1992) is a composite software quality metric that integrates HalsteadVolume, which quantifies the amount of information a reader must process to comprehend the code’s functionality, along with McCabe complexity and Size. The MI is computed using the following formula:
Other. The number of parameters (Parameters), the number of local variables (LocalVariables), and the Halstead Length (Length), which measures the total number of operators and operands, are also included alongside the other metrics. We also intended to include the FanIn metric in our analysis, but were unable to do so due to the complexity involved. For example, consider the Hadoop project, which contains approximately 70,000 methods. Since our goal is to make predictions at method inception, we need to capture code metrics at the time each method was first introduced. Measuring the FanIn metric would require constructing a separate call graph for the whole codebase at each method’s introduction commit. Unfortunately, building such call graphs is extremely time-consuming, as it requires preprocessing the entire codebase. Given that we work with 98 projects—each potentially requiring thousands of unique call graphs corresponding to thousands of distinct commits—this approach was computationally infeasible within the scope of our study. This limitation was mentioned in other similar studies as well (Chowdhury et al., 2024, 2025).